OpenStudy (anonymous):

Check if my answer is correct? True or false: i^44 + i^150 - i^74 - i^109 = -1 I think it's false, I got 1 - i.

OpenStudy (cwrw238):

you need to know that i^2 = -1 , i^4 = 1, i^3 = -1

OpenStudy (cwrw238):

sorry that last one i^3 = -i

OpenStudy (cwrw238):

so (i^4)^11 = I^44 = 1

OpenStudy (cwrw238):

i^150 = (i^4)^37 8 t^2 = -1

OpenStudy (cwrw238):

i^74 = (i^4)^18 * i^2 = -1 i^109 = (i^4)^27 * i = i

OpenStudy (cwrw238):

so we have 1 + (-1) - (-1) - i = 1 - i

OpenStudy (anonymous):

aah so i was right and it's false? :D

OpenStudy (cwrw238):

yes i agree with you

OpenStudy (anonymous):

thank you <3